### v1

# Purified pseudomode model for nonlinear system-bath interactions

Abstract
The theory of purified pseudomodes [arXiv:2412.04264 (2024)] was recently developed to provide a numerical tool for the analysis of the properties of a quantum system and the environment it couples to via linear system-bath interactions. Here we extend this theory to allow for the description of general nonlinear system-bath interactions. We demonstrate the validity of our method by considering the spontaneous decay of a two-level atom placed inside a single-mode lossy cavity and furthermore, its potential application to nanophotonics by calculating the resonance fluorescence spectrum of a quantum dot in the presence of a phonon environment. Our method provides a useful tool for the study of phonon-assisted emission in quantum dots and holds the the promise for broad applications in fields like quantum biology, nonlinear phononics, and nanophotonics.
I introduction
The manipulation and synthesis of photon- and phonon-matter interactions have achieved considerable progresses in past decades, making some challenging regimes experimentally accessible. Representative examples include, e.g., ultra- and deep-strong light-matter interactions in cavity and circuit quantum electrodynamics (QED) [1, 2, 3], time retardation effects caused by light propagation in waveguide QED [4, 5, 6] and giant atoms [7, 8, 9, 10], nonlinear optomechanical interactions induced by radiation pressure [11, 12, 13, 14], and nonlinear phononics [15, 16, 17], to name a few. A faithful and accurate characterization of the properties of physical systems in said regimes requires the improvement of existing or the invention of new analytical and numerical methods [18, 19].
The common starting point of devising such methods is often a microscopic description of the environment and its interaction with the system of interest, and a variety of perturbative or non-perturbative treatments can be derived depending on the parameter regime of interest. For example, perturbation theory can lead to an effective description of the system alone in terms of quantum master equations [18, 19, 20, 21, 22] in which the environmental effects are implicitly accounted for. On the other hand, non-perturbative methods can be grounded on a variety of ideas, ranging from discretization of the the continous environment [23, 24, 25] or the Feymann-Vernon influence functional [26, 27, 28, 29, 30, 31, 32, 33] for tensor-network simulations to solving unitarily equivalent models obtained from the reaction coordinate mapping [34, 35, 36, 37, 38, 39], the chain mapping [40, 41, 42, 43] and the polaron transformation [44, 45, 46, 47, 48, 49, 50], or constructing effective models with auxiliary degrees of freedom, e.g., the hierarchical equations of motion (HEOM) [51, 52, 53, 54, 55, 56, 57, 58, 59, 60], dissipatons [61, 62, 63, 64], and pseudomodes [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78].
In a previous paper [79], based on the pseudomode theory [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], some of us introduced a new class of ancillary damping modes and detailed a rigorous scheme to construct effective models allowing for practical simulations of the dynamics of a system and the input-output properties of its environment. These ancillary modes were named “purified pseudomodes” because their states effectively remain pure during the evolution, in stark contrast to pseudomodes which require a mixed-state description. This method is numerically exact and valid for thermal bosonic environments and for generic linear system-bath interactions, and only requires the spectral decompositions of the bath two-time correlation matrix.
In this paper, we aim to derive an extended purified pseudomode model for the simulation of open quantum systems in the presence of general nonlinear system-bath interactions, where the bath coupling operator is an analytical function of a linear coupling operator $X$, see the definition in Eq. (2). This topic remains largely unexplored in the literature, except for Refs. [80, 81] where the specific case of a system-bath interaction written as the sum of a linear term and a quadratic correction was considered in the context of the HEOM.
The analysis of nonlinear system-bath interactions, including the specific ones employed in Refs. [80, 81], is more challenging than the linear case because of the additional terms emerging in the Wick factorizations of multi-time bath correlation functions which depend on non-trivial equal-time contractions. Here, we show that the technical difficulties involved in re-summing all these extra Wick contractions due to nonlinearities are bypassed using the pseudomode theory. In fact, by its very definition, this method offers a natural one-to-one correspondence of multi-time correlation functions between the original bath and the corresponding pseudomode model, making the latter able to encode any nonlinear statistics of the original bath by mimicking the statistical properties on the linear operator $X$. We provide an easily implementable scheme to define the required pseudomode model and show that a further purification of the degrees of freedom of this model allows to optimize the overall memory requirements. The validity of our method is demonstrated for the problem of the spontaneous decay of a two-level atom strongly coupled to a single-mode lossy cavity. We also illustrate the potential applications of our method to more complex examples by studying the effects of nonlinearly-coupled phonons on the fluorescence spectrum of a two-level quantum dot.
The paper is organized as follows. In Sec. II, we describe the microscopic model of an open quantum system in which a system couples to a bosonic environment with nonlinear interactions, and show that extra bath statistics is required to reproduce the reduced state of the system. In Sec. III, we present the construction of an explicit purified pseudomode model allowing for the simulation of the reduced system dynamics. In Sec. IV, we numerically demonstrate the validity and applications of our method with two examples. A summary of the main results and an outlook are given in Sec. V.
II model and reduced system dynamics
We start by defining the general model of an open quantum system consisting of a system $S$ and a bosonic environment $B$, described by the Hamiltonian
$H=H_{S}+H_{B}+H_{\text{int}},$ (1)
where $H_{S}$ is the system Hamiltonian and $H_{B}=\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}$ is the bath Hamiltonian with $b_{k}$ ($b_{k}^{\dagger}$) denoting the annihilation (creation) operator of the bath harmonic mode having frequency $\omega_{k}$, satisfying $[b_{k},b_{k^{\prime}}^{\dagger}]=\delta_{kk^{\prime}}$. We consider the following general form for the system-bath interaction
$H_{\text{int}}=sQ(X)=s\sum_{n=1}^{\infty}\alpha_{n}X^{n},$ (2)
where $s$ and $Q$ are the system and bath coupling operators respectively, and $X=\sum_{k}(g_{k}b_{k}^{\dagger}+g_{k}^{*}b_{k})$ is a linear bath coupling operator with $g_{k}$ the coupling amplitudes. Here we have assumed that the bath coupling operator $Q$ is an analytical function of $X$, in other words, $Q$ can be expanded as a power series of $X$ with real-valued coefficients $\alpha_{n}\in\mathbb{R}$, so that $Q$ is Hermitian. In particular, the case $\alpha_{1}=1$ and $\alpha_{n\geq 2}=0$ corresponds to the linear system-bath interaction considered in most literature. In Refs. [80, 81], Xu and collaborators analyzed the case $\alpha_{n\geq 3}=0$ in the context of the HEOM.
We are interested in the reduced dynamics of the system $S$, encoded in the reduced density operator $\rho_{S}(t)\equiv\text{Tr}_{B}\rho(t)$, which is obtained by tracing over the environmental degrees of freedom in the full density matrix $\rho(t)$. The evolution of $\rho(t)$ obeys the von-Neumann equation $d\rho(t)/dt=-i[H,\rho(t)]$ in the Schrödinger picture. We also assume a factorizing initial state $\rho(0)=\rho_{S}(0)\otimes\rho_{\text{th}}$, where $\rho_{\text{th}}=\exp(-\beta H_{B})/\text{Tr}_{B}\exp(-\beta H_{B})$ denotes the bath thermal state at inverse temperature $\beta$.
In the interaction picture with respect to the free Hamiltonian $H_{0}=H_{S}+H_{B}$, the reduced system dynamics is described by
$\frac{d\rho_{S}^{I}(t)}{dt}=-i\text{Tr}_{B}[H_{\text{int}}^{I}(t),\rho_{S}^{I}(t)],$ (3)
where $\rho_{S}^{I}(t)=e^{iH_{0}t}\rho_{S}(t)e^{-iH_{0}t}$ and $H_{\text{int}}^{I}(t)=e^{iH_{0}t}H_{\text{int}}e^{-iH_{0}t}$ denote the corresponding reduced density operator and system-bath interaction Hamiltonian in the interaction picture, respectively. The formal solution of Eq. (3), namely, the Dyson series can be written as [73]
$\rho_{S}^{I}(t)=\sum_{m=0}^{\infty}\frac{(-i)^{m}}{m!}\rho_{m}^{I}(t),$ (4)
where the $m$th-order contribution is
[…]
with $\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}\equiv\text{Tr}_{B}\Big{[}\mathcal{T}_{B}\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rho_{\text{th}}\Big{]}$ denoting the multi-time bath correlation functions and the indexes $i_{1},\cdots,i_{m}\in\{l,r\}$ labelling the left- or right-action of the superoperators $\mathcal{S}_{i_{j}}^{t_{j}}$ and $\mathcal{Q}_{t_{j}}^{i_{j}}$. In other words, they are defined as $\mathcal{S}_{t_{j}}^{l}[\cdot]=s(t_{j})[\cdot]$, $\mathcal{S}_{t_{j}}^{r}[\cdot]=-[\cdot]s(t_{j})$, $\mathcal{Q}_{t_{j}}^{l}[\cdot]=Q(t_{j})[\cdot]$, and $\mathcal{Q}_{t_{j}}^{r}[\cdot]=[\cdot]Q(t_{j})$ where $s(t_{j})=\exp(iH_{B}t_{j})s\exp(-iH_{B}t_{j})$ and $Q(t_{j})=\exp(iH_{B}t_{j})Q\exp(-iH_{B}t_{j})$ correspond to the operators $s$ and $Q$ in the interaction picture, respectively. Note that, the minus sign in the definition of $\mathcal{S}_{t_{j}}^{r}$ is necessary to recover the commutators involving $H_{\text{int}}^{I}(t)$ in the Dyson series in Eq. (4). The time-ordering operators $\mathcal{T}_{S}$ and $\mathcal{T}_{B}$ ensure that the superoperators $\mathcal{S}_{i_{j}}^{t_{j}}$ and $\mathcal{Q}_{t_{j}}^{i_{j}}$ are in chronological order, respectively.
In the case of linear system-bath interaction $Q=X$, the superoperator version of Wick’s theorem proved in Ref. [73] allows to conclude that each multi-time correlation function $\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ appearing in Eq. (5) can be written as a sum of total contraction of the linear operator $X$. More precisely, each multi-time correlation function $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ can be reduced to a product of two-time correlators which are related to the bath two-time correlation function
[…]
where $t=t_{2}-t_{1}$ denote the time difference and $J(\omega)\equiv\pi\sum_{k}\lvert g_{k}\rvert^{2}\delta(\omega-\omega_{k})$ is the spectral density. The time-translational invariance of $C(t)$ is a consequence of the stationarity of the bath initial state $\rho_{\text{th}}$ under the bath free dynamics, manifested in the commutation relation $[H_{B},\rho_{\text{th}}]=0$. In addition, $C(t)$ satisfies the time-reversal property $C(t)=C^{*}(-t)$ with the symbol * representing the complex conjugation. Ultimately, the formal series in Eq. (4) can be re-summed to yield a compact expression $\rho_{S}^{I}(t)=\exp(-\mathcal{F}(t,s,C(t))\rho_{S}(0)$ where the influence superoperator $\mathcal{F}(t,s,C(t)$ is explicitly written as [73]
$\begin{array}[]{lll}\mathcal{F}[\boldsymbol{\cdot}]&\equiv&\displaystyle\int_{0}^{t}dt_{2}\int_{0}^{t_{2}}dt_{1}\Big{(}C(t_{1}-t_{2})[\boldsymbol{\cdot}s(t_{1}),s(t_{2})]\\ &&\displaystyle~{}~{}~{}-C(t_{2}-t_{1})[s(t_{1})\boldsymbol{\cdot},s(t_{2})]\Big{)}.\end{array}$ (7)
This compact expression, or equivalently its path integral representation, constitutes the starting point of several nonperturbative methods [26, 27, 28, 30, 29, 31, 32, 33, 52, 54, 51, 59, 58, 53, 55, 60, 56, 57, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78] developed for linear system-bath interaction.
While no compact expression can be found for $\rho_{S}^{I}(t)$ in the general case of nonlinear system-bath interactions considered in Eq. (2), Wick’s theorem [73] still allows to conclude that each multi-time correlation function $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{th}$ in Eq. (5) can be written as a sum of total contraction of the linear operator $X$. In this case, the resulting factorizations of $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ acquire explicit dependencies on the second moment of $X$, i.e.
$\sigma^{2}\equiv\text{Tr}_{B}\left[X(t_{1})X(t_{1})\rho_{\text{th}}\right]=C(0).$ (8)
As an example, the factorization of $\langle\mathcal{Q}_{t_{2}}^{i_{2}}\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ reads
$\begin{array}[]{lll}\langle\mathcal{Q}_{t_{2}}^{i_{2}}\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}&\displaystyle=\langle\mathcal{X}_{t_{2}}^{i_{2}}\mathcal{X}_{t_{2}}^{i_{2}}\rangle_{\text{th}}\langle\mathcal{X}_{t_{1}}^{i_{1}}\mathcal{X}_{t_{1}}^{i_{1}}\rangle_{\text{th}}\\ &\displaystyle~{}~{}~{}+2\langle\mathcal{X}_{t_{2}}^{i_{2}}\mathcal{X}_{t_{1}}^{i_{1}}\rangle_{\text{th}}\langle\mathcal{X}_{t_{2}}^{i_{2}}\mathcal{X}_{t_{1}}^{i_{1}}\rangle_{\text{th}},\end{array}$ (9)
where the moment $\sigma^{2}$ and the two-time correlation function $C(t)$ appear in the first and second lines, respectively. This explicit dependence on the moment $\sigma^{2}$ in the Wick factorizations of $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$, has critical and profound consequences on the construction of the purified pseudomode model, constituting a major challenge which we overcome in the next section.
III Purified pseudomode model for nonlinear system-bath interactions
In this section, we explicitly construct a purified pseudomode model that allows to reproduce the reduced system dynamics of the nonlinear open-system model described in Sec. II. To this end, we first present a brief summary of the conventional pseudomode theory in Sec. III.1. We then establish an explicit pseudomode model that reproduces both the two-time correlation (6) and the second moment (8) in Sec. III.2, based on spectral decompositions of the two-time correlation (6). In Sec. III.3, we generalize the purification technique in our previous work [79] to reach a purified pseudomode model for the special case $Q=X^{n}$ and ultimately, in Sec. III.4, we derive the desired purified pseudomode model for the general nonlinear system-bath interaction (2).
III.1 The pseudomode theory
In the conventional pseudomode theory [65, 68, 69], where often only linear system-bath interaction $Q=X$ is considered, the original continuum environment is modeled as an effective, discrete one so that they yield identical reduced dynamics of the system $S$. The effective environment consists of a number of discrete ancillary bosonic modes, each subject to an additional Markovian dissipation channel, so that the state $\rho_{\text{pm}}(t)$ describing the original system $S$ and the effective environment obeys the following Lindblad master equation in the Schrödinger picture
$\begin{array}[]{lll}\displaystyle\frac{d\rho_{\text{pm}}}{dt}&\displaystyle=-i[H_{\text{S}}+\sum_{\alpha}\Omega_{\alpha}d_{\alpha}^{\dagger}d_{\alpha}+sQ_{\text{pm}},\rho_{\text{pm}}]\\ &\displaystyle~{}~{}~{}+\sum_{\alpha}\Gamma_{\alpha}(2d_{\alpha}\rho_{\text{pm}}d_{\alpha}^{\dagger}-d_{\alpha}^{\dagger}d_{\alpha}\rho_{\text{pm}}-\rho_{\text{pm}}d_{\alpha}^{\dagger}d_{\alpha}),\end{array}$ (10)
where the coupling operator $Q_{\text{pm}}=X_{\text{pm}}$ is related to the linear coupling operator
$X_{\text{pm}}\equiv\sum_{\alpha}X_{\text{pm},\alpha}=\sum_{\alpha}\lambda_{\alpha}(d_{\alpha}+d_{\alpha}^{\dagger}),$ (11)
and $d_{\alpha}$ ($d_{\alpha}^{\dagger}$) denote the annihilation (creation) operator of the $\alpha$th pseudomode with frequency $\Omega_{\alpha}$ and decay rate $\Gamma_{\alpha}$. In Eq. (10), we have used a simplified version of the general pseudomode model [68, 69, 73], namely, each pseudomode is in contact with its own background environment held at zero temperature, resulting in a single Lindblad dissipator with decay rate $\Gamma_{\alpha}$ in the second line of Eq. (10). For consistency, this also implies that the initial state of the pseudomode model in Eq. (10) should be $\rho_{\text{pm}}(0)=\rho_{S}\otimes\rho_{\text{vac}}$ with $\rho_{\text{vac}}=\otimes_{\alpha}|0_{\alpha}\rangle\langle 0_{\alpha}|$ the vacuum state of all pseudomodes. Note that this zero-temperature assumption, alongside the independent pseudomode assumption underlying Eq. (10), can both be relaxed to further enlarge predictivity of the pseudomode model [69, 70, 73, 72, 74]. For example, all model parameters can be analytically continued to the complex plane such that they become complex-valued, i.e., $\Omega_{\alpha},\Gamma_{\alpha},\lambda_{\alpha}\in\mathbb{C}$ [69, 73]. On the other hand, mode-mode interactions can also be introduced to form a pseudomode network [70, 72, 74].
The equivalence between the original environment $B$ and the effective one made of pseudomodes is guaranteed if the latter can reproduce (at least approximately) the original bath correlation function $C(t)$ (6), i.e.,
$C(t)\cong C_{\text{pm}}(t),$ (12)
where $C_{\text{pm}}(t)\equiv\text{Tr}_{\text{pm}}[X_{\text{pm}}(t)X_{\text{pm}}\rho_{\text{vac}}]$ can be solved analytically as [73]
$C_{\text{pm}}(t)=\sum_{\alpha}\lambda_{\alpha}^{2}e^{-i\Omega_{\alpha}t-\Gamma_{\alpha}\lvert t\rvert}.$ (13)
In other words, for linear system-bath interaction, this equivalence condition ensures that the reduced density matrix $\rho_{S}(t)$ for the open-system model in Sec. II can also be obtained by tracing out in $\rho_{\text{pm}}(t)$ all pseudomode degrees of freedom, i.e., $\rho_{S}(t)=\text{Tr}_{\text{pm}}\rho_{\text{pm}}(t)$.
The above consideration can be easily generalized to nonlinear system-bath interactions (2). This is achieved by simply substituting in Eq. (10)
$Q_{\text{pm}}=\sum_{n=1}^{\infty}\alpha_{n}X_{\text{pm}}^{n},$ (14)
and again by satisfying the equivalence condition (12).
It is important to note that the equivalence in Eq. (12) is not required to be point-like. In fact, $C(t)$ always appears in the integrand of the integral in Eq. (5), meaning that the difference $\lvert C(t)-C_{\text{pm}}(t)\rvert$ can even be allowed to be arbitrarily large over a countable collection of times $\mathcal{S}=\{t|\lvert C(t)-C_{\text{pm}}(t)\rvert~{}\text{can be arbitrarily large}\}$. From this formal point of view, the explicit emergence of the moment in Eq. (8) in the Wick factorizations of the multi-time correlation function $\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ for the nonlinear interaction (2) requires a precise approximation of the correlation at $t=0$, i.e., $(t=0)\notin\mathcal{S}$, in stark constrast to the linear case. In Sec. III.2 we show that this crucial distinction leads to an extra zero-frequency pseudomode with respect to the linear case.
III.2 Explicit pseudomode model for the nonlinear system-bath interaction (2)
We now determine explicitly all unknown pseudomode parameters $\Omega_{\alpha}$, $\Gamma_{\alpha}$, $\lambda_{\alpha}$ in Eqs. (10) and (11), so that both the two-time correlation (6) and the second moment (8) are reproduced up to a controlled error.
As in the analysis of Ref. [79], we assume a spectral decomposition of its positive contribution $C(t\geq 0)$ in the form of
$C(t\geq 0)\cong\sum_{l=1}^{N}w_{l}e^{-(i\nu_{l}+\gamma_{l})t},$ (15)
with $w_{l}\in\mathbb{C}$, $\nu_{l}\in\mathbb{R}$, $\gamma_{l}\in\mathbb{R}_{+}$, and the condition $\sum_{l=1}^{N}w_{l}\cong\sigma^{2}$ is satisfied for consistency. Such an exponential decomposition can be realized using several existing algorithms, e.g., the AAA algorithm and the Prony decomposition, see Ref. [82] and the references therein for a comprehensive review. Because of the time-reversal property $C(t)=C^{*}(-t)$, the decomposition in Eq. (15) also gives the decomposition of the negative contribution $C(t\leq 0)$, written as
$C(t\leq 0)\cong\sum_{l=1}^{N}w_{l}^{*}e^{-(i\nu_{l}-\gamma_{l})t},$ (16)
with $\sum_{l}w_{l}^{*}\cong\sigma^{2}$ valid by construction.
To match the pseudomode correlation ansatz (13), we combine the two decompositions in Eqs. (15) and (16) to obtain the following approximation
$C(t)\cong C_{\text{pm}}^{a}(t)=C_{\text{pm},>}^{a}(t)+C_{\text{pm},<}^{a}(t)+C_{\text{zf}}(t),$ (17)
in the limit $a\to+\infty$, with $C_{\text{zf}}(t)=-\sigma^{2}e^{-a\lvert t\rvert}$ and
$C_{\text{pm},\gtrless}^{a}(t)=\sum_{l=1}^{N}(\text{Re}\,w_{l}\pm i\text{Im}\,w_{l})e^{-i(\nu_{l}\pm ia)t-(\gamma_{l}+a)\lvert t\rvert},\$ (18)
where $a\in[0,+\infty)$ is a free parameter. The use of the symbol $\cong$ in Eq. (17) highlights that the approximating error is dominated by the one made using the approximation ansatz in Eq. (15) and it requires further justification, which we now provide. Comparing Eq. (18) with Eq. (15) and Eq. (16), we find that
$\begin{array}[]{lll}C_{\text{pm},\gtrless}^{a}(t\gtrless 0)&\cong&C(t\gtrless 0),\\ C_{\text{pm},\gtrless}^{a}(t\lessgtr 0)&\sim&\mathcal{O}(e^{-a\lvert t\rvert})\;,\end{array}$ (19)
defining the intended forward- and backward-time nature of, respectively, $C^{a}_{\text{pm},>}(t)$ and $C^{a}_{\text{pm},<}(t)$, in the limit $a\to+\infty$. In other words, for $t\neq 0$, the overall approximation error $\lvert C(t)-C_{\text{pm}}^{a}(t)|$ arising from the introducing of the free parameter $a$ is of order $\mathcal{O}(e^{-a\lvert t\rvert})$, which vanishes in the limit $a\to+\infty$, justifying $C(t)\cong\lim_{a\to+\infty}C_{\text{pm}}^{a}(t)$ for $t\neq 0$. Further analysis for the case $t=0$ is necessary due to the explicit dependence of the dynamics on $C(0)$ in the nonlinear case analyzed here, as shown at the end of section Sec. III.1. Specifically, we introduced the additional term $C_{\text{zf}}(t)$ in Eq. (17) to correctly reproduce the second moment of $X(0)$, as can be explicitly checked as
$C_{\text{pm}}^{a}(0)=\sum_{l}(w_{l}+w_{l}^{*})-\sigma^{2}\cong 2\sigma^{2}-\sigma^{2}=\sigma^{2}=C(0)\;.$ (20)
This allows to conclude that, also at $t=0$, the error $\lvert C(t)-C_{\text{pm}}^{a}(t)|$ made in using Eq. (17) is dominated by the spectral decomposition in Eq. (15), which in principle can be made arbitrarily small. This leads to conclude that $C(t)\cong\lim_{a\to+\infty}C_{\text{pm}}^{a}(t)$ is valid for all $t\in\mathbb{R}$, ultimately justifying Eq. (17).
We can now safely proceed by comparing the expressions of $C_{\text{pm},\gtrless}^{a}(t)$ and $C_{\text{zf}}(t)$ with the pseudomode correlation ansatz (13), to conclude that each exponential on the right-hand side of Eq. (18) can be related to a zero-temperature pseudomode labeled by the multi-index $\alpha=(l\pm)$ and specified by the parameters
$\Omega_{l\pm}=\nu_{l}\pm ia,~{}\Gamma_{l\pm}=\gamma_{l}+a,~{}\lambda_{l\pm}=\sqrt{\text{Re}w_{l}\pm i\text{Im}\,w_{l}}.$ (21)
Similarly, $C_{\text{zf}}(t)$ corresponds to one zero-frequency pseudomode characterized by the parameters
$\Omega_{0}=0,~{}\Gamma_{0}=a,~{}\lambda_{0}=i\sigma.$ (22)
In this way, we have thus defined a pseudomode model consisting of $12N+$ 1 zero-temperature pseudomodes, so that the sum of their free correlations $C_{\text{pm}}^{a}(t)$ approximates $C(t)$ with the error dominating by the spectral decomposition in Eq. (15) in the limit $a\to+\infty$. This completes the construction of an explicit pseudomode model for general nonlinear system-bath interactions (2).
III.3 Purified pseudomode model for the special case $Q=X^{n}$
In our previous work [79], we have shown that when the frequency $\Omega$ and decay rate $\Gamma$ of a zero-temperature pseudomode is analytically continued along either of the paths $\Omega\to\Omega\pm ia$ and $\Gamma\to\Gamma+a$, the state of the pseudomode becomes effectively pure when the limit $a\to+\infty$ is reached, leading to an optimization of the pseudomode Hilbert space which can be exploited in numerical simulations. This important result allows us to introduce the concept of purified pseudomodes, which, formally, can be understood as pseudomodes characterizied by the frequency $\Omega\pm i\infty$ and the decay rate $\Gamma+\infty$. Interestingly, the purified pseudomode correlations, obtained by taking the limit $C_{\text{ppm}}^{\pm}(t)=\lim_{a\to+\infty}\lambda^{2}\exp(-i(\Omega\pm ia)t-(\Gamma+a)\lvert t\rvert)=\lambda^{2}\exp(-i\Omega t\mp\Gamma t)\Theta(\pm t)$, correspond to either the positive-time ($+$) or negative-time (--) contributions of the pseudomode correlation, see Eq. (13).
In Appendix. III.3, we generalize the purification technique in the previous work [79] to the special case $Q=X^{n}$, and show that in the limit $a\to+\infty$ the equation of motion in Eq. (10) becomes the following Lindblad-like master equation
$i\frac{d\rho_{\text{ppm}}(t)}{dt}=[H_{S},\rho_{\text{ppm}}]+\mathcal{L}_{0}\rho_{\text{ppm}}+\mathcal{L}_{\text{int},n}\rho_{\text{ppm}},$ (23)
with the generators $\mathcal{L}_{0}$ and $\mathcal{L}_{\text{int},n}$ defined as
$\begin{array}[]{lll}&\displaystyle\mathcal{L}_{0}[\boldsymbol{\cdot}]=\sum_{l=1}^{N}\Big{[}(\nu_{l}-i\gamma_{l})d_{l+}^{\dagger}d_{l+}[\boldsymbol{\cdot}]-(\nu_{l}+i\gamma_{l})[\boldsymbol{\cdot}]d_{l-}^{\dagger}d_{l-}\Big{]},\\ &\displaystyle\mathcal{L}_{\text{int},n}[\boldsymbol{\cdot}]=X_{+}^{n}s[\boldsymbol{\cdot}]-[\boldsymbol{\cdot}]sX_{-}^{n}+s[\boldsymbol{\cdot}]X_{-,d^{\dagger}}^{n}-X_{+,d}^{n}[\boldsymbol{\cdot}]s\\ &\displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{r=1}^{n-1}\binom{n}{r}\bigg{[}X_{+}^{n-r}s[\boldsymbol{\cdot}]X_{-,d^{\dagger}}^{r}-X_{+,d}^{r}[\boldsymbol{\cdot}]sX_{-}^{n-r}\bigg{]},\end{array}$ (24)
where $\rho_{\text{ppm}}(t)$ denotes the density matrix describing the state of the system $S$ and all purified pseudomodes, and we have defined the operators $X_{\pm}=X_{\pm,d}+X_{\pm,d^{\dagger}}$ with $X_{\pm,d}=\sum_{l=1}^{N}\lambda_{l\pm}d_{l\pm}$ and $X_{\pm,d^{\dagger}}=\sum_{l=1}^{N}\lambda_{l\pm}d_{l\pm}^{\dagger}$ their components to simplify the notation. The initial state is still $\rho_{\text{ppm}}(0)=\rho_{S}\otimes\rho_{\text{vac}}$. The generator (24) exhibits the common feature of purified pseudomodes, namely, the annihilation and creation operators of any purified pseudomode always appear on the same side of the density matrix. Specifically, $d_{l+}$ and $d_{l+}^{\dagger}$ ($d_{l-}$ and $d_{l-}^{\dagger}$) only act on the left-hand (right-hand) side in Eq. (24). Since all purified pseudomodes evolve from their vacuum, this critical feature directly implies that the states of all purified pseudomodes can effectively be described in terms of pure states during the evolution.
As a closing remark of this subsection, we note that by taking the limit $a\to+\infty$, the error related to the free parameters $a$ in the approximation (17) vanishes so that the only error source of the purified pseudomode model in Eq. (23) comes from the spectral decompositions in Eqs. (15) and (16). In addition, the equation of motion in Eq. (23) can be recast in the form of the HEOM, as shown explicitly in Appendix. B.
III.4 Purified pseudomode model for the nonlinear system-bath interaction (2)
The purification procedure described in Appendix. III.3 is additive, in the sense that for the system-bath interaction $H_{\text{int}}=sX^{n}+sX^{m}$ the generator $\mathcal{L}_{\text{ppf}}$ of the corresponding purified pseudomode model is simply $\mathcal{L}_{\text{ppf}}=\mathcal{L}_{0}+\mathcal{L}_{\text{int},n}+\mathcal{L}_{\text{int},m}$. This fact means that, for the general nonlinear interaction (2), the corresponding generator can be directly written as $\mathcal{L}_{\text{ppf}}=\mathcal{L}_{0}+\sum_{n=1}^{\infty}\alpha_{n}\mathcal{L}_{\text{int},n}$. This concludes our construction of the purified pseudomode model.
Figure 1: Spontaneous decay of a two-level atom placed inside a lossy cavity. Colorful solid curves are obtained from the corresponding purified pseudomode model (abbreviated as “PPM”) for four different combinations of the coefficients $(\alpha_{1},\alpha_{2},\alpha_{3})$, while black dashed lines are the results of the master equation in Eq. (25). Simulation parameters are $\omega_{S}=0.5$, $\nu=0.5$, $\gamma=0.1$, and $\lambda=0.3$.
[…]
IV Numerical Demonstrations
We proceed to illustrate the validity and potential application of our method with two examples. The numerical results presented in this section are obtained using QuTiP master equation solvers in the excitation-number restricted basis [83, 84, 85].
IV.1 Spontaneous decay of a two-level atom in a lossy cavity
For benchmarking, we first consider the spontaneous decay of a two-level atom strongly coupled to a lossy cavity with the resonance frequency $\nu$ and decay rate $\gamma$, so that the state $\rho_{\text{ac}}(t)$ describing both the atom and the cavity obeys the following Lindblad master equation
$\begin{array}[]{lll}\displaystyle\frac{d\rho_{\text{ac}}(t)}{dt}&=&\displaystyle-i[H_{S}+\nu b^{\dagger}b+sQ_{\text{ac}},\rho_{\text{ac}}]\\ &&\displaystyle+\gamma(2b\rho_{\text{ac}}b^{\dagger}-b^{\dagger}b\rho_{\text{ac}}-\rho_{\text{ac}}b^{\dagger}b),\end{array}$ (25)
where $s=\sigma_{x}$, $H_{S}=\omega_{S}\sigma_{z}/2$ is the bare atomic Hamiltonian with $\omega_{S}$ the level splitting, $b$ and $b^{\dagger}$ denote the annihilation and creation operators of the cavity, respectively. We consider a generalized light-matter interaction consisting of both the linear term and the second- and third-order corrections. In other words, the cavity coupling operator $Q_{\text{ac}}$ is written as $Q_{\text{ac}}=\alpha_{1}X_{\text{c}}+\alpha_{2}X_{\text{c}}^{2}+\alpha_{3}X_{\text{c}}^{3}$ with $X_{\text{c}}=\lambda(b+b^{\dagger})$ and $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathbb{R}$. The lossy cavity can be viewed as the environment of the two-level atom and its free correlation can be solved analytically as $C(t)=\lambda^{2}\exp(-i\nu t-\gamma\lvert t\rvert)$. The model considered here can be engineered in superconducting quantum circuits [86, 87]. Since its bath correlation function is just a single exponential, no extra spectral decomposition is needed, which we take advantage of to demonstrate the exactness of the purification procedure described in Sec. III.3.
In Fig. 1, we plot the dynamics of the atomic population and atom-cavity entanglement measured by the von-Neumann entropy for four different groups of the coefficients $(\alpha_{1},\alpha_{2},\alpha_{3})$ for a resonant cavity $\nu=\omega_{S}$. We compare the results obtained from the purified pseudomode model (solid lines) with those from the master equation in Eq. (25) (dashed black lines) and find full agreement in all cases. The presence of the small quadratic correction with $\alpha_{2}=0.4$ slightly increases the hybridization in the steady-state, while the the entanglement between the atom and the cavity decreases, as indicated by the red lines in Fig. 1. On the other hand, by tuning the sign of the cubic correction, stronger ($\alpha_{2}=0.15$, green lines) or weaker ($\alpha_{2}=-0.15$, purple lines) light-matter hybridization and entanglement can be engineered, possibly offering an extra control knob for quantum information processing and dissipative state engineering [88]. Note that, we have checked that the sign of the quadratic correction has no influence on the hybridization and entanglement (not shown).
Despite the general validity of the purified pseudomode model constructed in Sec. III, its numerical simulations may encounter instability issues for large nonlinearity, which arises from the necessary truncation of the infinite-dimensional Hilbert spaces of purified pseudomodes. In fact, similar issues may also occur for pseudomode models involving unphysical model parameters and the HEOM, and several solutions to circumvent numerical instability have been proposed [89, 90] in the latter context. In our simulations, we have carefully chosen model parameters to avoid such instability. A more careful study of this practical issue will be left for future works.
IV.2 Fluorescence spectrum of quantum dots in the presence of phonon environments
We next consider a more complex physical example, namely, the fluorescence spectrum of a two-level quantum dot in the presence of a phonon environment, which affects the excitation energy of the quantum dot, leading to the bath coupling operator $s=\sigma_{+}\sigma_{-}$. To measure the fluorescence spectrum, the quantum dot is driven by a resonant laser so that in the frame rotating with the laser frequency the bare Hamiltonian of the quantum dot is $H_{S}=\Omega\sigma_{x}/2$ where $\Omega$ is the Rabi frequency. The quantum dot also exhibits a radiative decay with rate $\kappa$, corresponding to a Lindblad dissipator with the collapse operator $\sigma_{-}$. We employ the super-Ohmic spectral density $J(\omega)=\alpha_{p}\omega^{3}\exp(-\omega^{2}/2\omega_{b}^{2})$ for the phonon environment, where $\alpha_{p}$ measures the coupling strength and $\omega_{b}$ is the cutoff frequency. The spectrum is obtained from the Fourier transform of the two-time correlation function, i.e.,
[…]
where the subscript “ss” means the expectation is taken over the steady-state.
[…]
V Conclusions and outlook
In summary, we have presented a generalized purified pseudomode model capable of simulating the reduced system dynamics of open quantum systems in the presence of general system-bath interactions (2). We have pointed out that for nonlinear interactions, both the two-time bath correlation (6) involving the linear coupling operator $X$ and its second moment (8) emerge explicitly in the Wick factorizations of the multi-time correlation functions. We introduce an extra zero-frequency pseudomode to account for the second moment (8) and are able to generalize the purification technique developed in our previous work [79] to ultimately derive a purified pseudomode model for the general nonlinear system-bath interaction (2). We have demonstrated the validity and application of our method with two numerical examples. In particular, the capability of calculating the fluorescence spectrum of quantum dots nonlinearly coupled to phonon environments could facilitate the study of phonon-assisted emission in nanophotonics.
As an outlook, the results presented in this paper could be applied for the modeling of other types of nonlinear system-bath interactions, such as the ones in Refs. [86, 87] which can be engineered in waveguide QED. Another possible future direction is to further extend the purified pseudomode model here to allow for the computation of bath properties [79].
Acknowledgements.
We would like to thank Dr. W. Yang for interesting discussions about the role of nonlinearities in the pseudomode model. M.C. acknowledges support from NSFC (Grant No. 11935012) and NSAF (Grant No. U2330401). N.L. is supported by the RIKEN Incentive Research Program and by MEXT KAKENHI Grants No. JP24H00816 and No. JP24H00820.
Appendix A Detailed derivations for the purification of the pseudomode model
In this section, we show that the purification technique first presented in our previous work [79] can be generalized to the case $Q=X^{n}$, a special case of the general nonlinear interaction (2).
We start by rewriting the pseudomode model given by Eqs. (10), (14), (21) and (22) in the Liouville space
$\begin{array}[]{lll}\mathcal{L}&\displaystyle=\mathcal{H}_{S}\otimes\mathcal{H}_{S}\otimes\Big{(}\bigotimes_{l=1}^{N}\big{(}\mathcal{H}_{\text{pm},l+}\otimes\mathcal{H}_{\text{pm},l+}\big{)}\otimes\mathcal{H}_{0}\Big{)}\\ &\displaystyle~{}~{}~{}\otimes\Big{(}\bigotimes_{l=1}^{N}\big{(}\mathcal{H}_{\text{pm},l-}\otimes\mathcal{H}_{\text{pm},l-}\big{)}\otimes\mathcal{H}_{0}\Big{)},\end{array}$ (27)
where $\mathcal{H}_{S}$, $\mathcal{H}_{\text{pm},l\pm}$ and $\mathcal{H}_{0}$ denote the Hilbert space of the system, the pseudomode indexed by $\alpha=l\pm$ and the zero-frequency pseudomode, respectively. To achieve this, we employ the “bra-flipper” superoperator $\mathfrak{U}$ introduced in Ref. [94] to define the mapping
$\mathfrak{U}[|\psi\rangle\langle\psi^{\prime}|]=|\psi\rangle\otimes|\psi^{\prime}\rangle^{*}\equiv|\psi\psi^{\prime}\rangle\rangle,$ (28)
for two state vectors $|\psi\rangle$ and $|\psi^{\prime}\rangle$ in a Hilbert space. We denote orthonormal basis states of the system $S$ as $|s\rangle$ and $|s^{\prime}\rangle$, and denote Fock states corresponding to the pseudomode indexed by $\alpha=l\pm$ ($\alpha=0$) as $|n_{l\pm}\rangle$ and $|n_{l\pm}^{\prime}\rangle$ ($|n_{0}\rangle$ and $|n_{0}^{\prime}\rangle$). In this basis, the density matrix $\rho_{\text{pm}}$ in Eq.(10) can be expanded as
$\begin{array}[]{lll}&\displaystyle\rho_{\text{pm}}=\sum_{s,s^{\prime},n_{l\pm},n_{l\pm}^{\prime}}(\rho_{\text{pm}})_{sn_{l+}n_{l-},s^{\prime}n_{l+}^{\prime}n_{l-}^{\prime}}\times\\ &\displaystyle|s\rangle\langle s^{\prime}|\otimes\Big{(}\bigotimes_{l=1}^{N}\big{(}|n_{l+}\rangle\langle n_{l+}^{\prime}|\otimes|n_{l-}\rangle\langle n_{l-}^{\prime}|\big{)}\Big{)}\otimes|n_{0}\rangle\langle n_{0}^{\prime}|.\end{array}$ (29)
Applying the bra-flipper superoperator $\mathfrak{U}$, we can vectorize $\rho_{\text{pm}}$ as
$\begin{array}[]{lll}&\displaystyle|\rho_{\text{pm}}\rangle\rangle=\sum_{s,s^{\prime},n_{l\pm},n_{l\pm}^{\prime}}(\rho_{\text{pm}})_{sn_{l+}n_{l-},s^{\prime}n_{l+}^{\prime}n_{l-}^{\prime}}\times\\ &\displaystyle|ss^{\prime}\rangle\rangle\otimes\Big{(}\bigotimes_{l=1}^{N}\big{(}|n_{l+}n_{l+}^{\prime}\rangle\rangle\otimes|n_{l-}n_{l-}^{\prime}\rangle\rangle\big{)}\Big{)}\otimes|n_{0}n_{0}^{\prime}\rangle\rangle.\end{array}$ (30)
On the other hand, this mapping of the density matrix also gives rise to the mapping of the left (right) action of an operator $O$, with support on either $\mathcal{H}_{\text{S}}$ or $\mathcal{H}_{\text{PM},l\pm}$ or $\mathcal{H}_{0}$, to $\bar{O}\equiv O\otimes I$ ($\tilde{O}\equiv I\otimes O^{\intercal}$), with support on either $\mathcal{H}_{\text{S}}\otimes\mathcal{H}_{\text{S}}$ or $\mathcal{H}_{\text{pm},l\pm}\otimes\mathcal{H}_{\text{pm},l\pm}$ or $\mathcal{H}_{0}\otimes\mathcal{H}_{0}$. Here, $I$ denotes the corresponding identity operator and $O^{\intercal}$ the transpose of $O$.
Ultimately, in the Liouville space, the equation of motion in Eq. (10) can be rewritten as
$\begin{array}[]{lll}\displaystyle i\frac{d|\rho_{\text{pm}}^{a}\rangle\rangle}{dt}=\mathfrak{L}|\rho_{\text{pm}}^{a}\rangle\rangle=(\mathfrak{L}_{S}+\mathfrak{L}_{\text{pm},0}^{a}+\mathfrak{L}_{\text{PM},\text{int}}^{a})|\rho_{\text{pm}}^{a}\rangle\rangle,\end{array}$ (31)
with $\mathfrak{L}_{\text{S}}=\bar{H}_{\text{S}}-\tilde{H}_{\text{S}}^{\intercal}$ and
$\begin{array}[]{lll}\displaystyle\mathfrak{L}_{\text{pm},0}^{a}&\displaystyle=\sum_{l=1}^{N}\Big{[}(\nu_{l}-i\gamma_{l})\bar{d}_{l+}^{\dagger}\bar{d}_{l+}-(\nu_{l}+i(\gamma_{l}+2a))\tilde{d}_{l+}^{\dagger}\tilde{d}_{l+}\\ &\displaystyle~{}~{}~{}+2i(\gamma_{l}+a)\bar{d}_{l+}\tilde{d}_{l+}+(\nu_{l}-i(\gamma_{l}+2a))\bar{d}_{l-}^{\dagger}\bar{d}_{l-}\\ &\displaystyle~{}~{}~{}-(\nu_{l}+i\gamma_{l})\tilde{d}_{l-}^{\dagger}\tilde{d}_{l-}+2i(\gamma_{l}+a)\bar{d}_{l-}\tilde{d}_{l-}\\ &\displaystyle~{}~{}~{}+2ia\bar{d}_{0}\tilde{d}_{0}-ia\bar{d}_{0}^{\dagger}\bar{d}_{0}-ia\tilde{d}_{0}^{\dagger}\tilde{d}_{0}\Big{]},\end{array}$ (32)
and
$\mathfrak{L}_{\text{PM},\text{int}}^{a}=\bar{s}\Big{[}\bar{X}_{+}+\bar{X}_{-}+\bar{X}_{0}\Big{]}^{n}-\tilde{s}^{\intercal}\Big{[}\tilde{X}_{+}+\tilde{X}_{-}+\tilde{X}_{0}\Big{]}^{n},$ (33)
where for later convenience we have introduced the operators $\bar{X}_{\xi}=\bar{X}_{\xi,d}+\bar{X}_{\xi,d^{\dagger}}$, $\tilde{X}_{\xi}=\tilde{X}_{\xi,d}+\tilde{X}_{\xi,d^{\dagger}}$ with $\xi\in\{+,-,0\}$, whose components are $\bar{X}_{\pm,d}=\sum_{l=1}^{N}\lambda_{l\pm}\bar{d}_{l\pm}$, $\bar{X}_{\pm,d^{\dagger}}=\sum_{l=1}^{N}\lambda_{l\pm}\bar{d}_{l\pm}^{\dagger}$, $\tilde{X}_{\pm,d}=\sum_{l=1}^{N}\lambda_{l\pm}\tilde{d}_{l\pm}$, $\tilde{X}_{\pm,d^{\dagger}}=\sum_{l=1}^{N}\lambda_{l\pm}\tilde{d}_{l\pm}^{\dagger}$, $\bar{X}_{0,d}=\lambda_{0}\bar{d}_{0}$, $\bar{X}_{0,d^{\dagger}}=\lambda_{0}\bar{d}_{0}^{\dagger}$, $\tilde{X}_{0,d}=\lambda_{0}\tilde{d}_{0}$ and $\tilde{X}_{0,d^{\dagger}}=\lambda_{0}\tilde{d}_{0}^{\dagger}$, respectively. In addition, the initial state $\rho_{\text{pm}}(0)$ is now mapped to the vector
$|\rho_{\text{pm}}(0)\rangle\rangle=|\rho_{\text{S}}(0)\rangle\rangle\otimes\bigotimes_{l=1}^{N}|\bar{0}_{l+}\tilde{0}_{l+}\bar{0}_{l-}\tilde{0}_{l-}\rangle\rangle\otimes|\bar{0}_{0}\tilde{0}_{0}\rangle\rangle,$ (34)
where the vacuum states of the modes $\bar{d}_{l\pm/0}$ and $\tilde{d}_{l\pm/0}$ are denoted as $|\bar{0}_{l\pm/0}\rangle$ and $|\tilde{0}_{l\pm/0}\rangle$, respectively.
To gain more intuition, we reiterate the argument in the previous work [79]. We observe that the decay rates of the modes $\tilde{d}_{l+}$, $\bar{d}_{l-}$, $\tilde{d}_{0}$ and $\bar{d}_{0}$ depend on the free parameter $a$, meaning that in the limit $a\to+\infty$ these modes decay so fast that they should be frozen in their initial vacuum states. This implies that in this limit these modes can be adiabatically eliminated so that they are dynamically decoupled from the other degrees of freedom.
To confirm this intuitive picture, we expand the interaction Liouville operator $\mathfrak{L}_{\text{PM},\text{int}}^{a}$ as
$\begin{array}[]{lll}\mathfrak{L}_{\text{PM},\text{int}}^{a}&\displaystyle=\bar{s}\sum_{r=0}^{n}\binom{n}{r}\bar{X}_{+}^{n-r}\bar{X}_{-0}^{r}-\tilde{s}^{\intercal}\sum_{r=0}^{n}\binom{n}{r}\tilde{X}_{+0}^{r}\tilde{X}_{-}^{n-r}\\ &\displaystyle=\bar{s}\sum_{r=0}^{n}\sum_{p=0}^{r}\binom{n}{r}\binom{r}{p}\bar{X}_{+}^{n-r}\bar{X}_{-0,d^{\dagger}}^{p}\bar{X}_{-0,d}^{r-p}\\ &\displaystyle~{}~{}~{}-\tilde{s}^{\intercal}\sum_{r=0}^{n}\sum_{p=0}^{r}\binom{n}{r}\binom{r}{p}\tilde{X}_{-}^{n-r}\tilde{X}_{+0,d^{\dagger}}^{p}\tilde{X}_{+0,d}^{r-p},\end{array}$ (35)
where in the first line we have introduced the shorthand notations $\bar{X}_{-0}=\bar{X}_{-}+\bar{X}_{0}$ and $\tilde{X}_{-0}=\tilde{X}_{-}+\tilde{X}_{0}$, and in the second line we have used the fact that the component $\bar{X}_{-0,d}$ ($\tilde{X}_{+0,d}$) commutes with $\bar{X}_{-0,d^{\dagger}}$ ($\tilde{X}_{+0,d^{\dagger}}$), i.e.,
$\begin{array}[]{lll}[\bar{X}_{-0,d},\bar{X}_{-0,d^{\dagger}}]&=\sum_{l=1}^{N}\lambda_{-}^{2}+\lambda_{0}^{2}=0,\\ {[\tilde{X}_{+0,d},\tilde{X}_{+0,d^{\dagger}}]}&=\sum_{l=1}^{N}\lambda_{+}^{2}+\lambda_{0}^{2}=0,\end{array}$ (36)
since the equalities $\sum_{l=1}^{N}\lambda_{l-}^{2}=\sum_{l=1}^{N}w_{l}^{*}=\sigma^{2}$, $\sum_{l=1}^{N}\lambda_{l+}^{2}=\sum_{l=1}^{N}w_{l}=\sigma^{2}$ and $\lambda_{0}^{2}=-\sigma^{2}$ are valid by construction, see Eqs. (15), (16) and (22). Based on Eq. (35), we further classify the terms in the interaction Liouville operator $\mathfrak{L}_{\text{PM},\text{int}}^{a}$ as
$\mathfrak{L}_{\text{PM},\text{int}}^{a}=\mathfrak{V}+\big{(}\bar{s}\bar{X}_{+}^{n}-\tilde{s}^{\intercal}\tilde{X}_{-}^{n}\big{)}+\mathfrak{R},$ (37)
with
$\begin{array}[]{lll}\displaystyle\mathfrak{V}&=&\displaystyle\bar{s}\sum_{r=1}^{n}\binom{n}{r}\bar{X}_{+}^{n-r}\bigg{[}\sum_{p=0}^{r}\binom{r}{p}\bar{X}_{-,d^{\dagger}}^{p}\bar{X}_{0,d^{\dagger}}^{r-p}\bigg{]}\\ &&\displaystyle-\tilde{s}^{\intercal}\sum_{r=1}^{n}\binom{n}{r}\tilde{X}_{-}^{n-r}\bigg{[}\sum_{p=0}^{r}\binom{r}{p}\tilde{X}_{+,d^{\dagger}}^{p}\tilde{X}_{0,d^{\dagger}}^{r-p}\bigg{]},\end{array}$ (38)
and
$\begin{array}[]{lll}\mathfrak{R}&\displaystyle=\bar{s}\sum_{r=1}^{n}\sum_{p=0}^{r-1}\binom{n}{r}\binom{r}{p}\bar{X}_{+}^{n-r}\bar{X}_{-0,d^{\dagger}}^{p}\bar{X}_{-0,d}^{r-p}\\ &\displaystyle~{}~{}~{}-\tilde{s}^{\intercal}\sum_{r=1}^{n}\sum_{p=0}^{r-1}\binom{n}{r}\binom{r}{p}\tilde{X}_{-}^{n-r}\tilde{X}_{+0,d^{\dagger}}^{p}\tilde{X}_{+0,d}^{r-p}.\end{array}$ (39)
The reason underlying this classification is that for the modes with enhanced dissipation, namely, $\tilde{d}_{l+}$, $\bar{d}_{l-}$, $\tilde{d}_{0}$ and $\bar{d}_{0}$, the term $\mathfrak{V}$ in Eq. (38) only involves their creation operators $\bar{X}_{-,d^{\dagger}}$, $\tilde{X}_{+,d^{\dagger}}$, $\bar{X}_{0,d^{\dagger}}$ and $\tilde{X}_{0,d^{\dagger}}$, meaning that $\mathfrak{V}$ can trigger transitions out of the vacuum of $\tilde{d}_{l+}$, $\bar{d}_{l-}$, $\tilde{d}_{0}$ and $\bar{d}_{0}$. Hence, the term $\mathfrak{V}$ should be eliminated to justify the aforementioned intuitive argument. In contrast, the term $\mathfrak{R}$ in Eq. (39) contains operator strings ending with either $\bar{X}_{-0,d}$ or $\tilde{X}_{+0,d}$. If $\mathfrak{V}$ can be adiabatically eliminated, the term $\mathfrak{R}$ has no influence on the dynamics, therefore can be simply dropped from the generator $\mathfrak{L}$.
We can now generalize the technique presented in the previous work [79] to adiabatically eliminate the term $\mathfrak{V}$ from the generator $\mathfrak{L}=\mathfrak{L}_{S}+\mathfrak{L}_{\text{pm},0}^{a}+\mathfrak{L}_{\text{PM},\text{int}}^{a}$. In particular, we look for a Liouvillian operator $\mathfrak{S}$, which, crucially, we require to be $\mathcal{O}(a^{-1})$, to define the Schrieffer-Wolff (SW) transformation
$e^{\mathfrak{S}}\mathfrak{L}e^{-\mathfrak{S}}=\mathfrak{L}+[\mathfrak{S},\mathfrak{L}]+\frac{1}{2!}[\mathfrak{S},[\mathfrak{S},\mathfrak{L}]]+\cdots,$ (40)
such that the following operator equation is satisfied
$[\mathfrak{S},\mathfrak{L}_{0}]=-\mathfrak{V}+\mathfrak{V}^{\prime}+\mathfrak{V}_{\text{res}}+\mathcal{O}(a^{-1}),$ (41)
with $\mathfrak{L}_{0}=\mathfrak{L}_{S}+\mathfrak{L}_{\text{pm},0}^{a}+\big{(}\bar{s}\bar{X}_{+}^{n}-\tilde{s}^{\intercal}\tilde{X}_{-}^{n}\big{)}+\mathfrak{R}$. Here, the operator $\mathfrak{V}^{\prime}$ represents the emerging mediated interactions, and for consistency we require, not to contain any operators related to the modes $\tilde{d}_{l+}$, $\bar{d}_{l-}$, $\tilde{d}_{0}$ and $\bar{d}_{0}$, and to be $\mathcal{O}(a^{0})$. The operator $\mathfrak{V}_{\text{res}}$ contains terms ending with either $\tilde{X}_{+,d}$ or $\bar{X}_{-,d}$ or $\tilde{X}_{0,d}$ or $\bar{X}_{0,d}$, which, for the same reason concerning the term $\mathfrak{R}$ in Eq.(37) renders $\mathfrak{V}_{\text{res}}$ irrelevant to the dynamics. Importantly, the existence of the commutator in Eq. (41) ensures nested commutators appearing in the SW transformation (40), such as $[\mathfrak{S},\cdots,[\mathfrak{S},\mathfrak{L}_{0}]]$ with at least two $\mathfrak{S}$, or $[\mathfrak{S},\cdots,[\mathfrak{S},\mathfrak{V}]]$ with at least one $\mathfrak{S}$, to be at least of order $\mathcal{O}(a^{-1})$. Hence, in the limit $a\to+\infty$, they vanish in the expansion of the SW transformation (40), leading to the simple generator
$e^{\mathfrak{S}}\mathfrak{L}e^{-\mathfrak{S}}=\mathfrak{L}_{0}+\mathfrak{V}^{\prime}+\mathfrak{V}_{\text{res}},$ (42)
under which the states of the modes $\tilde{d}_{l+}$, $\bar{d}_{l-}$, $\tilde{d}_{0}$ and $\bar{d}_{0}$ to be eliminated indeed remain on their initial vacuum states. Ultimately, this allows to discard $\mathfrak{R}$ and $\mathfrak{V}_{\text{res}}$ from Eq. (42), leading to an effective generator
$\begin{array}[]{lll}\displaystyle\mathfrak{L}_{n}&=&\displaystyle\mathfrak{L}_{S}+\sum_{l=1}^{N}\Big{[}(\nu_{l}-i\gamma_{l})\bar{d}_{l+}^{\dagger}\bar{d}_{l+}-(\nu_{l}+i\gamma_{l})\tilde{d}_{l-}^{\dagger}\tilde{d}_{l-}\Big{]}\\ &&\displaystyle+\bar{s}\bar{X}_{+}^{n}-\tilde{s}^{\intercal}\tilde{X}_{-}^{n}+\mathfrak{V}^{\prime}.\end{array}$ (43)
In addition, since $\mathfrak{S}$ is of order $\mathcal{O}(a^{-1})$, we have $\lim_{a\to+\infty}e^{\mathfrak{S}}|\rho_{\text{pm}}\rangle\rangle=|\rho_{\text{pm}}\rangle\rangle$, meaning that in the limit $a\to+\infty$ the vectorized density matrix remains invariant under the SW transformation.
The remaining task is to solve the operator equation in Eq. (41). We have found the solution
$\begin{array}[]{lll}&&\displaystyle-2ia\mathfrak{S}\\ &=&\displaystyle\bar{s}\Bigg{[}\sum_{p=0}^{n-1}\frac{1}{n-p}\bar{X}_{-,d^{\dagger}}^{n-p}\tilde{X}_{-,d}^{p}\\ &&\displaystyle~{}~{}~{}+\sum_{r=1}^{n-1}\sum_{p=0}^{r-1}\frac{1}{r-p}\binom{n}{r}\bar{X}_{-,d^{\dagger}}^{r-p}\tilde{X}_{-,d}^{p}\bar{X}_{+}^{n-r}\Bigg{]}\\ &&\displaystyle-\tilde{s}^{\intercal}\Bigg{[}\sum_{p=0}^{n-1}\frac{1}{n-p}\tilde{X}_{+,d^{\dagger}}^{n-p}\bar{X}_{+,d}^{p}\\ &&\displaystyle~{}~{}~{}~{}~{}~{}~{}+\sum_{r=1}^{n-1}\sum_{p=0}^{r-1}\frac{1}{r-p}\binom{n}{r}\tilde{X}_{+,d^{\dagger}}^{r-p}\bar{X}_{+,d}^{p}\tilde{X}_{-}^{n-r}\Bigg{]}\\ &&\displaystyle+\sum_{r=1}^{n}\frac{2}{r}\binom{n}{r}\bigg{[}\bar{s}\bar{X}_{+}^{n-r}\bar{X}_{0,d^{\dagger}}^{r}-\tilde{s}^{\intercal}\tilde{X}_{-}^{n-r}\tilde{X}_{0,d^{\dagger}}^{r}\bigg{]}\\ &&\displaystyle+\bar{s}\sum_{r=1}^{n}\binom{n}{r}\bar{X}_{+}^{n-r}\sum_{p=1}^{r-1}\binom{r}{p}\frac{2p}{r+p}\\ &&\displaystyle\times\bigg{[}\sum_{m=0}^{p-1}A_{m}\bar{X}_{-,d^{\dagger}}^{p-m}\bar{X}_{0,d^{\dagger}}^{r-p}\tilde{X}_{-,d}^{m}+\frac{2A_{p-1}}{r-p}\bar{X}_{0,d^{\dagger}}^{r-p}\tilde{X}_{-,d}^{p}\bigg{]}\\ &&\displaystyle-\tilde{s}^{\intercal}\sum_{r=2}^{n}\binom{n}{r}\tilde{X}_{-}^{n-r}\sum_{p=1}^{r-1}\binom{r}{p}\frac{2p}{r+p}\\ &&\displaystyle\times\bigg{[}\sum_{m=0}^{p-1}A_{m}\tilde{X}_{+,d^{\dagger}}^{p-m}\tilde{X}_{0,d^{\dagger}}^{r-p}\bar{X}_{+,d}^{m}+\frac{2A_{p-1}}{r-p}\tilde{X}_{0,d^{\dagger}}^{r-p}\bar{X}_{+,d}^{p}\bigg{]},\end{array}$ (44)
where the coefficients $A_{m}$, $m\in\mathbb{N}$, are defined as $A_{0}=1/p$ and for $m\geq 1$
$A_{m}=\frac{1}{p}\prod_{k=1}^{m}\frac{2(p-k+1)}{r+p-2k},$ (45)
satisfying the recurrence relation $A_{m}(r+p-2m)/2=A_{m-1}(p-m+1)$. Substituting the solution in Eq. (44) into Eq. (41) leads to the new mediated interaction
$\begin{array}[]{lll}\mathfrak{V}^{\prime}&=&\displaystyle\bar{s}\bigg{[}\tilde{X}_{-,d}^{n}+\sum_{r=1}^{n-1}\binom{n}{r}\bar{X}_{+}^{n-r}\tilde{X}_{-,d}^{r}\bigg{]}\\ &&\displaystyle-\tilde{s}^{\intercal}\bigg{[}\bar{X}_{+,d}^{n}+\sum_{r=1}^{n-1}\binom{n}{r}\bar{X}_{+,d}^{r}\tilde{X}_{-}^{n-r}\bigg{]},\\ \end{array}$ (46)
and the residual term
$\begin{array}[]{lll}&&\mathfrak{V}_{\text{res}}\\ &=&\displaystyle\bar{s}\sum_{r=1}^{n-1}\sum_{p=0}^{r-1}\frac{n-r}{r-p}\binom{n}{r}\bar{X}_{-,d^{\dagger}}^{r-p}\tilde{X}_{-,d}^{p}\bar{X}_{+}^{n-r-1}\tilde{X}_{+,d}\\ &&\displaystyle-\tilde{s}^{\intercal}\sum_{r=1}^{n-1}\sum_{p=0}^{r-1}\frac{n-r}{r-p}\binom{n}{r}\tilde{X}_{+,d^{\dagger}}^{r-p}\bar{X}_{+,d}^{p}\tilde{X}_{-}^{n-r-1}\bar{X}_{-,d}\\ &&\displaystyle+2\sum_{r=1}^{n}\binom{n}{r}\bigg{[}\bar{s}\bar{X}_{+}^{n-r}\bar{X}_{0,d^{\dagger}}^{r-1}\tilde{X}_{0,d}-\tilde{s}^{\intercal}\tilde{X}_{-}^{n-r}\tilde{X}_{0,d^{\dagger}}^{r-1}\bar{X}_{0,d}\bigg{]}\\ &&\displaystyle+\bar{s}\sum_{r=2}^{n}\binom{n}{r}\sum_{p=1}^{r-1}\binom{r}{p}\frac{2p}{r+p}\sum_{m=1}^{p-1}A_{m}\bar{X}_{-,d^{\dagger}}^{p-m}\tilde{X}_{-,d}^{m}\\ &&\displaystyle\times\bigg{[}(r-p)\bar{X}_{+}^{n-r}\bar{X}_{0,d^{\dagger}}^{r-p-1}\tilde{X}_{0,d}\\ &&\displaystyle~{}~{}~{}~{}+(n-r)\bar{X}_{+}^{n-r-1}\bar{X}_{0,d^{\dagger}}^{r-p}\tilde{X}_{+,d}\bigg{]}\\ &&\displaystyle-\tilde{s}^{\intercal}\sum_{r=2}^{n}\binom{n}{r}\sum_{p=1}^{r-1}\binom{r}{p}\frac{2p}{r+p}\sum_{m=0}^{p-1}A_{m}\tilde{X}_{+,d^{\dagger}}^{p-m}\bar{X}_{-,d}^{m}\\ &&\displaystyle\times\bigg{[}(r-p)\tilde{X}_{-}^{n-r}\tilde{X}_{0,d^{\dagger}}^{r-p-1}\tilde{X}_{0,d}\\ &&\displaystyle~{}~{}~{}~{}+(n-r)\tilde{X}_{-}^{n-r-1}\tilde{X}_{0,d^{\dagger}}^{r-p}\bar{X}_{-,d}\bigg{]}\\ &&\displaystyle+\bar{s}\sum_{r=2}^{n}\binom{n}{r}\sum_{p=1}^{r-1}\binom{r}{p}\frac{4p}{r+p}A_{p-1}\tilde{X}_{-,d}^{p}\\ &&\displaystyle\times\bigg{[}\bar{X}_{+}^{n-r}\bar{X}_{0,d^{\dagger}}^{r-p-1}\tilde{X}_{0,d}+\frac{n-r}{r-p}\bar{X}_{+}^{n-r-1}\bar{X}_{0,d^{\dagger}}^{r-p}\tilde{X}_{+,d}\bigg{]}\\ &&\displaystyle-\tilde{s}^{\intercal}\sum_{r=2}^{n}\binom{n}{r}\sum_{p=1}^{r-1}\binom{r}{p}\frac{4p}{r+p}A_{p-1}\bar{X}_{+,d}^{p}\\ &&\displaystyle\times\bigg{[}\tilde{X}_{-}^{n-r}\tilde{X}_{0,d^{\dagger}}^{r-p-1}\bar{X}_{0,d}+\frac{n-r}{r-p}\tilde{X}_{-}^{n-r-1}\tilde{X}_{0,d^{\dagger}}^{r-p}\bar{X}_{-,d}\bigg{]}.\end{array}$ (47)
Switching back to the density matrix representation, we arrive at the superoperators in Eq. (24).
Appendix B Recasting Eq. (23) into the hierarchical equations of motion
To rewrite Eq. (23) in the form of the HEOM, we denote Fock states of purified pseudomodes as $|\mathbf{m},\mathbf{n}\rangle$, where the sequences $\mathbf{m}=(m_{1},m_{2},\cdots)$ and $\mathbf{n}=(n_{1},n_{2},\cdots)$ specify the occupations $m_{l}$ and $n_{l}$ of the modes $d_{l\pm}$, respectively. As pointed out in the main text, in Eq. (24) the annihilation and creation operators of a purified pseudomode always appear on the same side of the density matrix. This observation combined with the assumption that all purified pseudomdoes are initialized in their vacuum allow us to conclude that only the matrices $\rho_{\mathbf{m},\mathbf{n}}\equiv\langle\mathbf{m},\mathbf{0}|\rho_{\infty}|\mathbf{0},\mathbf{n}\rangle$ are nonvanishing, which correspond to the auxiliary density matrices (ADOs) in the HEOM. These ADOs have the same dimensions as the reduced density matrix $\rho_{S}$. According to Eq. (23), the ADOs $\rho_{\mathbf{m},\mathbf{n}}$ obey the equations of motion
$\begin{array}[]{lll}&\displaystyle id\rho_{\mathbf{m},\mathbf{n}}/dt\\ &\displaystyle=[H_{\text{S}},\rho_{\mathbf{m},\mathbf{n}}]\\ &\displaystyle+\sum_{l}\Big{(}(\nu_{l}-i\gamma_{l})m_{l}\rho_{\mathbf{m},\mathbf{n}}-(\nu_{l}+i\gamma_{l})n_{l}\rho_{\mathbf{m},\mathbf{n}}\Big{)}\\ &\displaystyle+\sum_{\mathbf{m}^{\prime}}\langle\mathbf{m}|X_{+}^{n}|\mathbf{m}^{\prime}\rangle s\rho_{\mathbf{m}^{\prime},\mathbf{n}}-\sum_{\mathbf{n}^{\prime}}\langle\mathbf{n}^{\prime}|X_{-}^{n}|\mathbf{n}\rangle\rho_{\mathbf{m},\mathbf{n}^{\prime}}s\\ &\displaystyle+\sum_{\mathbf{n}^{\prime}}\langle\mathbf{n}^{\prime}|X_{-,d^{\dagger}}^{n}|\mathbf{n}\rangle s\rho_{\mathbf{m},\mathbf{n}^{\prime}}-\sum_{\mathbf{m}^{\prime}}\langle\mathbf{m}|X_{+,d}^{n}|\mathbf{m}^{\prime}\rangle\rho_{\mathbf{m}^{\prime},\mathbf{n}}s\\ &\displaystyle+\sum_{r=1}^{n-1}\binom{n}{r}\sum_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}\bigg{[}\langle\mathbf{m}|X_{+}^{n-r}|\mathbf{m}^{\prime}\rangle\langle\mathbf{n}^{\prime}|X_{-,d^{\dagger}}^{n-r}|\mathbf{n}\rangle s\rho_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}\\ &\displaystyle-\langle\mathbf{m}|X_{+,d}^{n-r}|\mathbf{m}^{\prime}\rangle\langle\mathbf{n}^{\prime}|X_{-}^{n-r}|\mathbf{n}\rangle\rho_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}s\bigg{]},\end{array}$ (48)
which take the explicit form of the HEOM. The reduced density matrix of the system $S$ corresponds to $\rho_{S}=\rho_{\mathbf{0},\mathbf{0}}$. Note that, in the case of linear system-bath interaction $n=1$, the equations in Eq. (48) reduce to the free-pole HEOM, cf., Eq. (5) in the supplemental material of Ref. [60].